OpenStudy (anonymous):

plzz I need help to find the surface are of a cone with radius 17 and slant height of 32 any idea some one very smart could help me plzzzz

OpenStudy (jdoe0001):

 so what do you think would be the height of the cone? keeping in mind that \(\bf \textit{Surface Area of a cone} = \boldsymbol{\pi r\sqrt{r^2+h^2}+\pi r^2}\\ \quad \\ h = height\\ r = radius\)